OpenStudy (anonymous):

A tugboat has completely filled it’s 2034 L fuel tank. If it uses 113L/h, how many liters remain after a 7 hour trip?

OpenStudy (texaschic101):

7 hr trip at 113L per hr = 7 * 113 = 791 L 2034 - 791 = 1243 L remaining

OpenStudy (anonymous):

@Dianegrace I like your profile picture ;p